Oxidation number of chromium in K2Cr2O7


1
Expert's answer
2021-03-15T09:44:54-0400

Oxidation number of K = +1

oxidation number of O = -2

Let the oxidation number of Cr =X

So

+1(2) + X(2) + -2(7) = 0

2 + X(2) - 14 = 0

-12 + X(2) = 0

2X = 12

X=6

So the oxidation number of Cr in K2Cr2O7 is +6